OpenStudy (anonymous):

8. One of the Roman dice in the British Museum has 6 square faces and 8 triangular faces. It is twice as likely to land on any given square face as any given triangular face. What is the probability that the face it lands on is triangular, when thrown?

OpenStudy (unklerhaukus):

the probability is the number of favorable outcome divided by the total number of outcomes

OpenStudy (anonymous):

Huh? yes but what is the answer to the question? or how would i work it out?

OpenStudy (anonymous):

Basics: The prob of landing on any particular square face is equal to the probability for any other square face. They are all equal. The prob of landing on any particular triangular face is equal to the probability for any other triangular face. They are all equal. OK with that?

OpenStudy (anonymous):

Let P(one particular triangular face) = p Then what is the total probability of landing on any one of the triangular faces?

OpenStudy (anonymous):

You there?

OpenStudy (anonymous):

sorry back mate, was having dinner

OpenStudy (anonymous):

back

OpenStudy (anonymous):

Total prob of landing on any tri face = 6p Prob of landing on any one square face = 2p Total prob of landing on any square face = 12p Total prob of landing on any face = 6p + 12p = is defined to be equal to 1

OpenStudy (anonymous):

Gotta go.

OpenStudy (anonymous):

okie doke cheers for the help mate ;)

OpenStudy (anonymous):

Thankyou very much

OpenStudy (anonymous):

the available answers were (A)4 (B) 3 (C)3 (D) 3 (E)2 7 11 7 10 5

OpenStudy (anonymous):

oh hold on one sec sorry wrong ones

OpenStudy (anonymous):

uhhh they are all fractions, cant post them

OpenStudy (anonymous):

And I messed up, 8 triangular faces Total prob of landing on any face = 8p + 12p = is defined to be equal to 1 p = 1/20 can you do it now?

OpenStudy (anonymous):

a) 4/7 b) 3/11, c) 3/7, d) 3/10, and e) 2/5

OpenStudy (anonymous):

uhh im having a go again now

OpenStudy (anonymous):

8 tri faces each with a p = 1/20 total p if it could be any tri face = 8/20

OpenStudy (anonymous):

yes that works, that means the answer is e 2/5

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

gosh i see how easy it is now, my calculus is easier than that... problem solving is not my strong point

OpenStudy (anonymous):

i really appreciate it

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

i have a nation wide maths competition tomorrow. I can do conventional maths like calculus/trig/ whatever else reasonably well, but i cannot problem solve, and all fo these questions are multi choice problemsolving